# Preserving Angles Improves Feature Distillation of Foundation Models

### v2

@@ 4 Experiments: Feature distillation
In this section, the CosPress feature distillation approach is compared to Proteus and distilled variants of the DINOv2 models. While we do not have access to the proprietary LVD-142M dataset used to distill the DINOv2 models, it has been shown that ImageNet-1K [36] is sufficient to distill models with comparable accuracy across a range of measures [56].
@@ 4.1 Experimental setup
Vision Transformer [11] models are distilled using larger DINOv2 teachers on the ImageNet-1K [36] training dataset, comprising 1000 categories across more than 1.2 million images. To enable a fair comparison, we reproduce the results of the Proteus paper and train CosPress models using a unified codebase. This ensures consistency in the optimizers, samplers, augmentations and other hyperparameters. Following the DINOv2 kNN evaluation [47] and linear probing approach [32] with an additional batchnorm layer [24], evaluations are undertaken on the ImageNet validation set, as well as nine fine-grained classification benchmarks (Oxford Pets [33], FGVC Aircraft [27], Describable Textures [9], Stanford Cars [21], CUB200 [46], CIFAR-10/100 [22], Flowers-102 [31] and Food-101 [7]) and the Pascal VOC 2012 segmentation task [12]. Performance is also tested on several robustness benchmarks including ImageNet-V2 [35], Sketch [43], ImageNet-R [15] and ImageNet-A [16].
We additionally consider the OpenOOD benchmarks [48]. Foundation models are trained on a diverse dataset and excel in this task, and whether distilled students can also perform well has not been previously considered. This section focuses on the ImageNet-1K OpenOOD benchmark, which uses SSB-hard [6] and NINCO [42] as near OOD data, and iNaturalist [19], OpenImage-O [44] and Describable Textures [9] as far OOD data.
Table 2: ImageNet classification. Comparison of performance on ImageNet-1K under kNN and linear probing evaluation approaches.
Method Arch Teacher kNN Linear
Proteus ViT-Ti/14 DINOv2 ViT-S/14 73.1 76.1
Proteus-$V_{k}D$ ViT-Ti/14 DINOv2 ViT-S/14 73.0 75.9
CosPress ViT-Ti/14 DINOv2 ViT-S/14 74.3 76.8
DINOv2 ViT-S/14 79.0 81.1
Proteus ViT-S/14 DINOv2 ViT-B/14 79.8 82.0
CosPress ViT-S/14 DINOv2 ViT-B/14 80.4 82.3
@@
Method Arch Teacher kNN
DINOv2 Backbone Stu. head Tea. head Teacher
Proteus ViT-Ti/14 ViT-S/14 73.1 73.5 79.0
Proteus-$V_{k}D$ ViT-Ti/14 ViT-S/14 73.0 73.3 79.0
CosPress ViT-Ti/14 ViT-S/14 74.3 78.8 79.0
Proteus ViT-S/14 ViT-B/14 79.8 80.0 82.1
CosPress ViT-S/14 ViT-B/14 80.4 82.1 82.1
@@
Figure 4: Visualising orthogonality. Kernel density estimate plots of the diagonal and non-diagonal elements for the scaled Gram matrices of the linear maps $\mathbf{W}$ in the teacher and student heads, drawn from CosPress and Proteus respectively. The dashed coloured lines represent $\mathbf{W}^{\top}\mathbf{W}/\alpha$ and the solid lines represent $\mathbf{W}\mathbf{W}^{\top}/\beta$, where $\alpha,\beta$ are defined as in Tab. 4. A perfectly orthogonal matrix $\mathbf{W}$ will have a Gram matrix with density on the black dashed vertical lines.
@@
Method Distance to the Identity Measures
$\|\mathbf{A}-\mathbf{I}_{D_{T}}\|_{F}$ $\|\mathbf{B}-\mathbf{I}_{D_{S}}\|_{F}$ $\text{Tr}\left(|\mathbf{A}-\mathbf{I}_{D_{T}}|\right)$ $\text{Tr}\left(|\mathbf{B}-\mathbf{I}_{D_{S}}|\right)$
Proteus 27.3 9.5 57.1 17.0
Proteus-$V_{k}D$ 24.9 7.7 152.7 8.2
CosPress 20.3 2.6 3.4 4.2
@@
Method Arch Teacher Dataset
C10 C100 Food CUB DTD Pets Cars Aircr Flowers Average
Proteus ViT-Ti/14 DINOv2 ViT-S/14 95.1 81.4 83.5 84.1 72.9 94.2 72.8 54.1 96.0 81.6
CosPress ViT-Ti/14 DINOv2 ViT-S/14 94.9 81.9 84.6 85.1 73.8 94.1 75.3 55.7 96.8 82.5
DINOv2 ViT-S/14 [32] 97.7 87.5 89.1 88.1 80.6 95.1 81.6 74.0 99.6 88.1
Proteus ViT-S/14 DINOv2 ViT-B/14 97.8 87.7 89.7 88.4 78.0 95.9 82.8 62.9 97.6 86.8
CosPress ViT-S/14 DINOv2 ViT-B/14 97.8 87.6 90.3 88.9 78.0 95.9 84.0 63.4 98.8 87.2
@@
Method Arch Teacher mIoU
Proteus ViT-Ti/14 DINOv2 ViT-S/14 70.5
Proteus w/o patch loss ViT-Ti/14 DINOv2 ViT-S/14 69.7
CosPress ViT-Ti/14 DINOv2 ViT-S/14 71.1
DINOv2 ViT-S/14 81.2
Proteus ViT-S/14 DINOv2 ViT-B/14 77.3
Proteus w/o patch loss ViT-S/14 DINOv2 ViT-B/14 77.1
CosPress ViT-S/14 DINOv2 ViT-B/14 77.9
@@
Method Arch Teacher Test Dataset
DINOv2 IN-V2 Sketch IN-R IN-A
Proteus ViT-Ti/14 ViT-S/14 64.3 25.5 37.8 11.4
CosPress ViT-Ti/14 ViT-S/14 64.9 27.9 40.7 13.2
ViT-S/14 [32] 70.9 41.2 53.7 33.5
Proteus ViT-S/14 ViT-B/14 72.2 38.4 50.0 29.6
CosPress ViT-S/14 ViT-B/14 72.5 40.4 52.3 31.5
@@
Method Arch Teacher GPUs GPU hours GPU memory
Proteus ViT-Ti/14 DINOv2 ViT-S/14 1 92 55GB
CosPress ViT-Ti/14 DINOv2 ViT-S/14 1 95 47GB
Proteus ViT-S/14 DINOv2 ViT-B/14 2 182  [^1] GB
CosPress ViT-S/14 DINOv2 ViT-B/14 2 154 81GB
@@ 4.2 Results
CosPress trains more competitive students.
Tab. 2 shows that CosPress trains students with better performance in comparison to Proteus [56], for both the linear probing and kNN evaluation methods. It is also found that the teacher head can project the embeddings from the teacher network into the latent space of the student with minimal loss of kNN accuracy, and that the token student head from Proteus has a higher kNN accuracy than the model backbone (Tab. 3). These observations confirm the motivations for CosPress—the Proteus student heads are not an uninformative mapping into a higher dimensional space, but are contaminated with information relevant for reproducing the teacher model. Further, a high-quality projection that compresses the teacher emebeddings into the latent space of the student—that preserves cosine similarity—can be learnt, and this provides more effective supervision.
In Tab. 2 we also consider Proteus-$V_{k}D$, where the projection matrices $\mathbf{W}$ in the Proteus student head are constrained to be right-orthogonal using the $V_{k}D$ approach [29]. This method builds a re-parametrisation map using skew symmetry and a matrix exponential approximation to construct $\mathbf{W}$ such that it must be right-orthogonal. Tab. 2 shows that in this context, this re-parametrisation slightly harms performance, and does not totally prevent contamination of the student head $g_{\phi}$.
It is also observed in Tab. 2 that the CosPress and Proteus approaches can outperform the distilled DINOv2 models of the same size. However, it is challenging to determine if these distillation approaches are more effective, as the DINOv2 models were trained on a larger proprietary dataset, of which the ImageNet-1K dataset was only a small subset.
As theorised in Lemma 1, it is found that CosPress learns a linear map $\mathbf{W}$ in the teacher head (Eq. 12) that is approximately left and right-orthogonal, up to a scaling factor. More concisely, we find that
where $\alpha,\beta$ are positive real numbers. Qualitatively, this can be seen in Fig. 4 where kernel density plots are shown of the elements from Gram matrices formed using the linear maps $\mathbf{W},\mathbf{W}^{\top}$. These projections are taken from the CosPress teacher head and the Proteus class token student head obtained while training the ViT-Ti/14 student network. The scaled CosPress Gram matrices are much closer to the identity matrix, and this is measured quantitatively using the Frobenius norm and trace in Tab. 4.
While the Proteus-$V_{k}D$ approach in Fig. 4 and Tab. 4 does learn a projection $\mathbf{W}$ that is approximately right-orthogonal, there is a large degree of error. This is due to the approximation of the matrix exponential that is used in the $V_{k}D$ method [29], and CosPress is able to learn a right-orthogonal projection matrix with less error.
CosPress improves performance on classification tasks.
Tab. 5 shows that the models distilled by CosPress have improved or similar performance over Proteus for all downstream fine-grained classification tasks. Competitive accuracy is also achieved with the distilled DINOv2 models of the same size, which CosPress outperforms on six of the nine datasets. Poorest performance is achieved on the FGVC-Aircraft dataset, which likely reflects differences in the training data used for distillation. The LVD-142M dataset used to train and distill the DINOv2 models contains a million images with high similarity to the FGVC-Aircraft dataset [32], whereas ImageNet only contains a single airliner class with approximately 1300 images.
CosPress improves segmentation performance.
Tab. 6 shows that CosPress also improves accuracy on downstream segmentation tasks in comparison to Proteus. CosPress does not include the masked patch loss objective, that we confirm improves the performance of Proteus on dense tasks, and incorporating it into CosPress may improve performance further. The DINOv2 distilled model outperforms CosPress in this case. Further pretraining with an increased image resolution was found to be key to improving the performance of the DINOv2 models on dense tasks [32], but is not undertaken in training the CosPress and Proteus student models.
@@
DeiT [40] ViT-Ti/16 RegNetY-16GF ImageNet 93.1 77.7 77.7 93.3 98.3 87.8 89.9 93.0
@@
Method Arch Parameters Dataset
C10 C100 Food Pets
NAT [25] MobileNetV2 [37] 4.5-9.0M 98.4 88.3 89.4 94.3
CeiT [51] CeiT-T 6.4M 98.5 88.4 93.8
CosPress ViT-Ti/14 5.5M 98.8 89.6 92.7 94.9
CosPress distills a more robust student model.
Tab. 7 shows that CosPress results in improved performance over Proteus across a range of ImageNet-1K robustness and generalisation benchmarks. The DINOv2 distilled model obtains better performance in this instance for all benchmarks except ImageNet-V2, but CosPress closes the gap between the ImageNet-1K and LVD-142M distilled models significantly.
CosPress reproduces the OOD detection performance of the teacher.
CosPress is faithful to the teacher networks when it comes to OOD detection performance, as shown in Tab. 1. Proteus performs very poorly on this benchmark, with worse performance observed for higher quality models. In contrast, CosPress is able to distill models that have strong OOD performance, even outperforming their DINOv2 counterparts.
CosPress does not require additional computational resources.
Tab. 8 provides timings and GPU memory usage for fitting the Proteus and CosPress models described in this section. Training time is similar for the ViT-Ti/14 and DINOv2 ViT-S/14 student-teacher pair, but CosPress is more efficient for larger models. This is due to the masked patch loss in Proteus, which requires that the student network is evaluated once on unmasked inputs, and a second time on masked inputs. As a result, training is faster for CosPress with larger students. CosPress is also slightly more memory efficient compared to Proteus, and further computational savings could be made by freezing the teacher head $h_{\phi}$ once a sufficiently high quality map has been learned.
5 Experiments: Specialist models
5.1 Experimental setup
The CosPress models distilled in the previous section are compared with models that have been further finetuned with CosPress — an additional pretraining step where distillation is undertaken on a smaller target dataset of interest. The strong DeiT [40] pretrained weights are also considered, which were trained on ImageNet-1K with a larger CNN network using class-based knowledge distillation [17].
We focus on a small set of tasks, including CIFAR-10/100 [22], Food-101 [7] and Oxford Pets [33]. Oxford Pets is a challenging dataset for training specialised ViT networks, as it only contains 3680 training images. The DINOv2 linear probe evaluation method is employed [32], as well as finetuning using the DeiT recipe [40]. When training models with this latter approach, the linear prediction head is trained before finetuning the backbone, to avoid distorting the pretrained features [23].
5.2 Results
CosPress finetuning improves the performance of specialist models.
Tab. 9 shows that CosPress finetuning improves downstream performance, even when the training datasets are quite small. For every dataset considered, this additional pretraining step improves linear probe evaluations with a frozen backbone by a significant margin (1-5%). These benefits remain under the strong DeiT training recipe, which further finetunes the model backbone. While CIFAR-10/100 and Food-101 have much stronger results under DeiT finetuning, we find that Oxford Pets has best performance with a linear probe evaluation after CosPress finetuning. This reflects the small size of the Oxford pets dataset, which makes training ViT networks challenging. Further, Tab. 9 highlights the strong performance of CosPress finetuning over the pretrained DeiT baseline, with large improvements using a linear probe, and more moderate improvements (excluding Oxford pets) under DeiT finetuning.
@@ 5.2 Results
A ViT-Tiny network finetuned with CosPress can have competitive accuracy compared to other approaches in the literature that have been highly optimised to perform well on specialist tasks with a small and efficient model. Tab. 10 shows that CosPress finetuning trains competitive networks in comparison to Neural Architecture Search (NAT) [25] and Convolution-enhanced image Transformers (CeiT) [51]. Feature distillation methods like CosPress are an additional approach, that could be used in conjunction with these techniques to build highly performant lightweight vision models.
We also find that CosPress improves the generalisability and OOD detection performance for these specialist models, as shown in Tab. S11 and Tab. S12 in the supporting information.

### v3

@@ 4 Experiments: Feature distillation
In this section, the CosPress feature distillation approach is compared to Proteus and distilled variants of the DINOv2 models. While we do not have access to the proprietary LVD-142M dataset used to distill the DINOv2 models, it has been shown that ImageNet-1K (ILSVRC15) is sufficient to distill models with comparable accuracy across a range of measures (zhang2024accessingvisionfoundationmodels).
@@ 4.1 Experimental setup
Vision Transformer (dosovitskiy2020image) models are distilled using larger DINOv2 teachers on the ImageNet-1K (ILSVRC15) training dataset, comprising 1000 categories across more than 1.2 million training images. To enable a fair comparison, we reproduce the results of the Proteus paper and train CosPress models using a unified codebase. This ensures consistency in the optimizers, samplers, augmentations and other hyperparameters. Following Proteus (zhang2024accessingvisionfoundationmodels), student networks are distilled for 300 epochs using a batch size of 1024, cosine learning rate decay with five warmup epochs (loshchilov2016sgdr), an AdamW optimizer (loshchilov2017decoupled), a repeated augmentation sampler with three views per image (fort2021drawing), and RandAugment (cubuk2020randaugment) image augmentations (rw2019timm). An ablation study on the hyperparameters introduced by CosPress is provided in Appendix˜C of the supporting information.
Following the DINOv2 kNN evaluation (wu2018unsupervised) and linear probing approach (oquab2023dinov2) with an additional batchnorm layer (lee2023rethinking), evaluations are undertaken on the ImageNet validation set, as well as nine fine-grained classification benchmarks (Oxford Pets (parkhi2012cats), FGVC Aircraft (maji2013fine), Describable Textures (cimpoi2013describing), Stanford Cars (krause20133d), CUB200 (welinder2010caltech), CIFAR-10/100 (krizhevsky2009learning), Flowers-102 (nilsback2008automated) and Food-101 (bossard2014food)) and the Pascal VOC 2012 segmentation task (pascal-voc-2012). Performance is also tested on several robustness and generalisation benchmarks including ImageNet-V2 (recht2019imagenet), Sketch (wang2019learning), ImageNet-R (hendrycks2021many) and ImageNet-A (hendrycks2021natural).
We additionally consider the OpenOOD benchmarks (yang2022openood). Foundation models are trained on a diverse dataset and excel in this task, and whether distilled students can reproduce this performance has not been previously considered. This section focuses on the ImageNet-1K OpenOOD benchmark, which uses SSB-hard (bitterwolf2023out) and NINCO (vaze2022openset) as near OOD data, and iNaturalist (vanhorn2018inaturalist), OpenImage-O (Wang_2022_CVPR) and Describable Textures (cimpoi2013describing) as far OOD data.
4.2 Results
Table 2: ImageNet classification. Comparison of performance on ImageNet-1K under kNN and linear probing evaluation approaches. We report the mean and standard deviation over four runs with different random seeds for the Proteus and CosPress ViT-Ti/14 models.
Method Arch Teacher kNN Linear Proteus ViT-Ti/14 DINOv2 ViT-S/14 $73.0\pm 0.1$ $76.1\pm 0.2$ Proteus-$V_{k}D$ ViT-Ti/14 DINOv2 ViT-S/14 73.0 75.9 CosPress ViT-Ti/14 DINOv2 ViT-S/14 $74.3\pm 0.1$ $76.6\pm 0.1$ DINOv2 ViT-S/14 79.0 81.1 Proteus ViT-S/14 DINOv2 ViT-B/14 79.8 82.0 CosPress ViT-S/14 DINOv2 ViT-B/14 80.4 82.3
@@ 4.2 Results
Method Arch Teacher kNN DINOv2 Backbone Stu. head Tea. head Teacher Proteus ViT-Ti/14 ViT-S/14 73.1 73.5 79.0 Proteus-$V_{k}D$ ViT-Ti/14 ViT-S/14 73.0 73.3 79.0 CosPress ViT-Ti/14 ViT-S/14 74.3 78.8 79.0 Proteus ViT-S/14 ViT-B/14 79.8 80.0 82.1 CosPress ViT-S/14 ViT-B/14 80.4 82.1 82.1
CosPress trains more competitive students.
Table˜2 shows that CosPress trains students with better performance in comparison to Proteus (zhang2024accessingvisionfoundationmodels), for both the linear probing and kNN evaluation methods. These improvements are statistically significant, taking into account the low variability observed across different random seeds. It is also found that the teacher head can project the embeddings from the teacher network into the latent space of the student with minimal loss of kNN accuracy, and that the token student head from Proteus has a higher kNN accuracy than the model backbone (Table˜3). These observations confirm the motivations for CosPress—the Proteus student heads are not an uninformative mapping into a higher dimensional space, but are contaminated with information relevant for reproducing the teacher model. Further, a high-quality projection that compresses the teacher emebeddings into the latent space of the student—that preserves cosine similarity—can be learnt, and this provides more effective supervision.
In Table˜2 we also consider Proteus-$V_{k}D$, where the projection matrices $\mathbf{W}$ in the Proteus student head are constrained to be right-orthogonal using the $V_{k}D$ approach (miles2024vkd). This method builds a re-parametrisation map using skew symmetry and a matrix exponential approximation to construct $\mathbf{W}$ such that it is approximately right-orthogonal. Table˜2 shows that in this context, this re-parametrisation does not significantly impact performance, and does not completely prevent contamination of the student head $g_{\phi}$.
It is also observed in Table˜2 that the CosPress and Proteus approaches can outperform the distilled DINOv2 models of the same size. However, it is challenging to determine if these distillation approaches are more effective, as the DINOv2 models were trained on a larger proprietary dataset, of which the ImageNet-1K dataset was only a small subset.
@@
Figure 3: Visualising orthogonality. Kernel density estimate plots of the diagonal and non-diagonal elements for the scaled Gram matrices of the linear maps $\mathbf{W}$ in the teacher and student heads, drawn from CosPress and Proteus respectively. The dashed coloured lines represent $\mathbf{W}^{\top}\mathbf{W}/\alpha$ and the solid lines represent $\mathbf{W}\mathbf{W}^{\top}/\beta$, where $\alpha,\beta$ are defined as in Table˜4. A perfectly orthogonal matrix $\mathbf{W}$ will have a Gram matrix with density on the black dashed vertical lines.
@@
Method Distance to the Identity Measures $\|\mathbf{A}-\mathbf{I}_{D_{T}}\|_{F}$ $\|\mathbf{B}-\mathbf{I}_{D_{S}}\|_{F}$ $\text{Tr}\left(|\mathbf{A}-\mathbf{I}_{D_{T}}|\right)$ $\text{Tr}\left(|\mathbf{B}-\mathbf{I}_{D_{S}}|\right)$ Proteus 27.3 9.5 57.1 17.0 Proteus-$V_{k}D$ 24.9 7.7 152.7 8.2 CosPress 20.3 2.6 3.4 4.2
As theorised in Lemma 1, it is found that CosPress learns a linear map $\mathbf{W}$ in the teacher head (Eq.˜13) that is approximately left and right-orthogonal, up to a scaling factor. More concisely, we find that
where $\alpha,\beta$ are positive real numbers. Qualitatively, this can be seen in Fig.˜3 where kernel density plots are shown of the elements from Gram matrices formed using the linear maps $\mathbf{W},\mathbf{W}^{\top}$. These projections are taken from the CosPress teacher head and the Proteus class token student head obtained while training the ViT-Ti/14 student network. The scaled CosPress Gram matrices are much closer to the identity matrix, and this is measured quantitatively using the Frobenius norm and trace in Table˜4.
While the Proteus-$V_{k}D$ approach in Fig.˜3 and Table˜4 does learn a projection $\mathbf{W}$ that is approximately right-orthogonal, there is a large degree of error. This is due to the approximation of the matrix exponential that is used in the $V_{k}D$ method (miles2024vkd), and CosPress is able to learn a right-orthogonal projection matrix with less error.
@@
Method Arch Teacher Dataset C10 C100 Food CUB DTD Pets Cars Aircr Flowers Average Proteus ViT-Ti/14 DINOv2 ViT-S/14 95.1 81.4 83.5 84.1 72.9 94.2 72.8 54.1 96.0 81.6 CosPress ViT-Ti/14 DINOv2 ViT-S/14 94.9 81.9 84.6 85.1 73.8 94.1 75.3 55.7 96.8 82.5 DINOv2 ViT-S/14 97.7 87.5 89.1 88.1 80.6 95.1 81.6 74.0 99.6 88.1 Proteus ViT-S/14 DINOv2 ViT-B/14 97.8 87.7 89.7 88.4 78.0 95.9 82.8 62.9 97.6 86.8 CosPress ViT-S/14 DINOv2 ViT-B/14 97.8 87.6 90.3 88.9 78.0 95.9 84.0 63.4 98.8 87.2
CosPress improves performance on classification tasks.
Table˜5 shows that the models distilled by CosPress have improved or similar performance over Proteus for all downstream fine-grained classification tasks. Competitive accuracy is also achieved with the distilled DINOv2 models of the same size, which CosPress outperforms on six of the nine datasets. Poorest performance is achieved on the FGVC-Aircraft dataset, which likely reflects differences in the training data used for distillation. The LVD-142M dataset used to train and distill the DINOv2 models contains a million images with high similarity to the FGVC-Aircraft dataset (oquab2023dinov2), whereas ImageNet only contains a single airliner class with approximately 1300 images.
@@
Method Arch Teacher mIoU Proteus ViT-Ti/14 DINOv2 ViT-S/14 70.5 Proteus w/o patch loss ViT-Ti/14 DINOv2 ViT-S/14 69.7 CosPress ViT-Ti/14 DINOv2 ViT-S/14 71.1 DINOv2 ViT-S/14 81.2 Proteus ViT-S/14 DINOv2 ViT-B/14 77.3 Proteus w/o patch loss ViT-S/14 DINOv2 ViT-B/14 77.1 CosPress ViT-S/14 DINOv2 ViT-B/14 77.9
CosPress improves segmentation performance.
Table˜6 shows that CosPress also improves accuracy on downstream segmentation tasks in comparison to Proteus. CosPress does not include the masked patch loss objective, that we confirm improves the performance of Proteus on dense tasks, and incorporating it into CosPress may improve performance further. The DINOv2 distilled model outperforms CosPress in this case. Further pretraining with an increased image resolution was found to be key to improving the performance of the DINOv2 models on dense tasks (oquab2023dinov2), but is not undertaken in training the CosPress and Proteus student models.
@@
Method Arch Teacher Test Dataset DINOv2 IN-V2 Sketch IN-R IN-A Proteus ViT-Ti/14 ViT-S/14 64.3 25.5 37.8 11.4 CosPress ViT-Ti/14 ViT-S/14 64.9 27.9 40.7 13.2 ViT-S/14 (oquab2023dinov2) 70.9 41.2 53.7 33.5 Proteus ViT-S/14 ViT-B/14 72.2 38.4 50.0 29.6 CosPress ViT-S/14 ViT-B/14 72.5 40.4 52.3 31.5
CosPress distills a more robust student model.
Table˜7 shows that CosPress results in improved performance over Proteus across a range of ImageNet-1K robustness and generalisation benchmarks. The DINOv2 distilled model obtains better performance in this instance for all benchmarks except ImageNet-V2, but CosPress closes the gap between the ImageNet-1K and LVD-142M distilled models significantly.
CosPress reproduces the OOD detection performance of the teacher.
CosPress is faithful to the teacher networks when it comes to OOD detection performance, as shown in Table˜1. Proteus performs very poorly on this benchmark, with worse performance observed for larger student models. In contrast, CosPress is able to distill models that have strong OOD performance, even outperforming their DINOv2 counterparts.
Table 8: Feature distillation with different teachers. Comparison of performance on ImageNet-1K under kNN and linear probing evaluation methods with other kinds of teacher backbones, using different architectures and training approaches.
Method Arch Teacher kNN Linear Proteus ViT-T/14 DINOv2 ViT-B/14 w/reg 71.1 75.1 CosPress ViT-T/14 DINOv2 ViT-B/14 w/reg 74.0 76.4 Proteus ViT-T/16 CLIP ViT-B/16 63.6 71.4 CosPress ViT-T/16 CLIP ViT-B/16 64.0 72.0
CosPress improves performance across other teacher networks.
Table˜7 demonstrates that CosPress also trains higher performing student networks than Proteus when using CLIP (radford2021learning) and DINOv2 w/reg (darcet2023vision) teacher networks. These experiments employ the same hyperparameters as those in Table˜2. Additional results exploring feature distillation with ViT-T students and larger teacher networks are provided in Appendix˜B of the supporting information.
@@
Method Arch Teacher GPUs GPU hours GPU memory Proteus ViT-Ti/14 DINOv2 ViT-S/14 1 92 55GB CosPress ViT-Ti/14 DINOv2 ViT-S/14 1 95 47GB Proteus ViT-S/14 DINOv2 ViT-B/14 2 182  [^1] GB CosPress ViT-S/14 DINOv2 ViT-B/14 2 154 81GB
CosPress does not require additional computational resources.
Table˜9 provides timings and GPU memory usage for fitting the Proteus and CosPress models described in this section. Training time is similar for the ViT-Ti/14 and DINOv2 ViT-S/14 student-teacher pair, but CosPress is more efficient for larger models. This is due to the masked patch loss in Proteus, which requires that the student network is evaluated once on unmasked inputs, and a second time on masked inputs. As a result, training is faster for CosPress with larger students. CosPress is also slightly more memory efficient compared to Proteus, and further computational savings could be made by freezing the teacher head $h_{\phi}$ once a sufficiently high quality map has been learned.
5 Experiments: Specialist models
5.1 Experimental setup
The CosPress models distilled in the previous section are compared with models that have been further finetuned with CosPress—an additional pretraining step where distillation is undertaken on a smaller target dataset of interest. The strong DeiT (touvron2021training) pretrained weights are also considered, which were distilled from ImageNet-1K with a larger CNN network using class-based knowledge distillation (hinton2015distilling). The same hyperparameters and training methodology is used as in Section˜4, with the exception of the number of training and warmup epochs.
This section focuses on a set of small-scale tasks, including CIFAR-10/100 (krizhevsky2009learning), Food-101 (bossard2014food) and Oxford Pets (parkhi2012cats). We employ 300 training epochs and 10 warmup epochs for CIFAR-10/100, and 3000 training epochs and 100 warmup epochs for Oxford Pets. The DINOv2 linear probe evaluation method is employed (oquab2023dinov2), as well as finetuning using the DeiT recipe (touvron2021training). When training models with this latter approach, the linear prediction head is trained before finetuning the backbone, to avoid distorting the pretrained features (kumar2022fine).
@@ 5.2 Results
DeiT (touvron2021training) ViT-Ti/16 RegNetY-16GF ImageNet 93.1 77.7 77.7 93.3 98.3 87.8 89.9 93.0
@@
CosPress finetuning improves the performance of specialist models.
Table˜10 shows that CosPress finetuning improves downstream performance, even when the training datasets are quite small. For every dataset considered, this additional pretraining step improves linear probe evaluations with a frozen backbone by a significant margin (1-5%). These benefits remain under the strong DeiT training recipe, which further finetunes the model backbone. While CIFAR-10/100 and Food-101 have much stronger results under DeiT finetuning, we find that Oxford Pets has best performance with a linear probe evaluation after CosPress finetuning. This reflects the small size of the Oxford pets dataset, which makes training ViT networks challenging.
@@
Method Architecture Parameters Dataset C10 C100 Food Pets NAT (lu2021neural) MobileNetV2 (sandler2018mobilenetv2) 4.5-9.0M 98.4 88.3 89.4 94.3 CeiT (yuan2021incorporating) CeiT-T 6.4M 98.5 88.4 93.8 CosPress ViT-Ti/14 5.5M 98.8 89.6 92.7 94.9
@@
A ViT-Tiny network finetuned with CosPress can have competitive accuracy compared to other approaches in the literature that have been highly optimised to perform well on specialist tasks with a small and efficient model. Table˜11 shows that CosPress finetuning trains competitive networks in comparison to Neural Architecture Transfer (NAT) (lu2021neural) and Convolution-enhanced image Transformers (CeiT) (yuan2021incorporating). Feature distillation methods like CosPress are an additional approach, that could be used in conjunction with these techniques to build highly performant lightweight vision models.
Table 12: Specialist models — generalisability. Comparison of generalisability of specialist models on the cartoon subsets of the CIFAR-10-W benchmark (sun2024cifarwarehouse). We report mean per-class accuracy due to dataset imbalances. In-distribution training images (top) and cartoon images (bottom) are included for reference.
Method Arch Teacher Pretraining dataset Linear DeiT
Diff Bin Bai 360 Diff Bin Bai 360
DeiT (touvron2021training) ViT-Ti/16 RegNetY-16GF ImageNet 65.9 51.0 47.6 48.8 86.9 62.6 56.0 60.1
The challenging cartoon subsets of the CIFAR-10-W benchmark (sun2024cifarwarehouse) are used to test generalisation performance on CIFAR-10. Table˜12 shows that CosPress finetuning leads to improved generalisability for specialist models on CIFAR-10. Under a linear probing evaluation, CosPress finetuning strongly improves generalisability on two of the four datasets, and improves generalisability for all datasets even after DeiT finetuning.
Table 13: Specialist models — OOD detection. Comparison of performance on the OpenOOD benchmark (yang2022openood). The AUC is reported for detecting OOD images.
CIFAR-10 CIFAR-100 CIFAR-10 CIFAR-100
DeiT (touvron2021training) ViT-Ti/16 RegNetY-16GF ImageNet 57.01 47.04 58.14 46.19 96.79 98.69 87.45 86.73
The CIFAR-10/100 OpenOOD benchmarks (yang2022openood) are used to test OOD detection performance. Table˜13 shows that CosPress finetuning leads to improved performance on the OpenOOD (yang2022openood) benchmark for specalist models on CIFAR-10/100. Without DeiT finetuning, strong improvements in OOD detection are observed under CosPress finetuning over the ImageNet pretrained baselines. Some improvements remain after DeiT finetuning, but are smaller.
6 Discussion
Significance of angles in deep learning.
Language models have been shown to exhibit a principle of superposition, wherein concepts are encoded along nearly orthogonal directions in representation space (bricken2023monosemanticity). While only $d$ vectors can be exactly orthogonal in a $d$-dimensional space, high-dimensional geometry allows for the construction of up to $\exp(d)$ approximately orthogonal vectors (with pairwise cosine similarity less than $\epsilon>0$) enabling the representation of a vastly larger set of concepts in practice (elhage2022superposition). This phenomenon is closely related to the Johnson-Lindenstrauss lemma (freksen2021introduction), and similar properties have been observed in foundation models for computer vision (bhalla2024interpreting). By preserving angular relationships between image embeddings, CosPress maintains the semantic structure of the foundation model feature space in the student networks it trains.
Limitations.
Even with CosPress, a small generalisation gap remains. Models trained with CosPress do not generalise quite as well as the original DINOv2 distilled variants (Table˜7). Without access to the proprietary LVD-142M dataset, it is difficult to determine whether this gap arises from the limitation of performing feature distillation solely on ImageNet-1K, or from a shortcoming in the methodology itself.